y=-x+7, y-x=3
find out if the equation lines are perpendicular

Answer by chriswen(106)

so first you convert the equations to y=mx+b.
y=-x+7 ... (i)
y-x=3
y=x+3 ... (ii)
m is the slope.
(i) slope of -1
(ii) slope of 1
For two lines to be perpendicular the products of the slopes is -1.
m1 * m2=-1
(-1)*(1)=
-1=-1
Since the products of both slopes is -1 therefore both lines are perpendicular to each other.